# Sequentially Auditing Differential Privacy

### v1

@@ 1 Introduction
Our specific test statistics are built upon the Maximum Mean Discrepancy (MMD) gretton2012kernel , a powerful kernel-based metric for comparing probability distributions. MMD is particularly well-suited for the two-sample tests emerging in privacy auditing, as bounds on MMD can be easily translated to approximate DP parameters 2024DPAuditorium . Further, MMD is highly flexible. It can incorporate prior knowledge in white-box settings —such as access to intermediate gradients or the noise distribution family— or, as our experiments show, operate effectively in a black-box manner.
@@
  • We introduce a new general MMD-based one-sided sequential testing framework. As a key application, we instantiate this framework for auditing approximate DP (see Section˜3). Our method enables anytime valid inference while controlling Type I error, eliminating the need for pre-defined test sample sizes, and ensuring a bounded expected stopping time under the alternative. We establish these theoretical guarantees in Theorems 3.2 and 3.3.
@@
  • We validate our methods on common DP mechanisms with Gaussian and Laplace noise. We further demonstrate efficacy on auditing benchmark algorithms (2024DPAuditorium, ; dpsniper, ) and provide results for the challenging case of DP-SGD abadi2016deep ; song2013stochastic , showcasing the practical benefits of early failure detection enabled by our sequential approach.
@@ 3 Sequential Auditing
This section formalizes the proposed auditing by betting framework. We start by formally introducing the problem statement and continue by introducing an MMD-based test statistic. Subsequently, we develop Algorithm 1, a sequential test for privacy auditing. We then provide formal guarantees on controlling Type I error uniformly over the null, and demonstrating exponential growth under the alternative, which implies a finite time stopping time when detecting faulty algorithms. We conclude this section with a comparative discussion of two subroutines within the main sequential algorithm.
@@ 3.1 Problem statement
Theorem˜3.1 offers a uniform improvement over of (2024DPAuditorium, , Theorem 4.13) and provides a strict improvement for any $\varepsilon>0$. Further, the new bound has the nice property of not becoming vacuous as $\varepsilon$ increases. In fact, it approaches $\sqrt{2}$ as $\varepsilon\to\infty$, whereas the previous bound grows to infinity with $e^{\varepsilon}$.
@@ 3.2 An abstract template for $(\varepsilon,\delta)$-DP sequential testing
$\mathcal{K}_{0}^{*}=1,\quad\mathcal{K}_{t}^{*}(\lambda)=\mathcal{K}^{*}_{t-1}(\lambda)\times(1+\lambda[f^{*}(X_{t})-f^{*}(Y_{t})-\tau(\varepsilon,\delta)]),$ (3)
@@
We are not the first to draw inspiration from shekhar2023nonparametric2sampletesting to design a one-sided test. Recently, richter2024auditing proposed a one-sided test in the context of detecting behavioral shifts in language models for the hypotheses $H_{0}:D(\mathbb{P},\mathbb{Q})\leq\tau$ versus $H_{1}:D(\mathbb{P},\mathbb{Q})>\tau$ for some $\tau>0$, where the distance function is defined as $D(\mathbb{P},\mathbb{Q})=\sup_{f\in\mathcal{F}}\left|\mathbb{E}_{X\sim\mathbb{P},\,Y\sim\mathbb{Q}}[f(X)-f(Y)]\right|$ and $\mathcal{F}$ is a class of neural networks with scalar outputs in [1/2,1/2][-1/2,1/2]. Their method also constructs a stochastic process that is a supermartingale under $H_{0}$ and can grow under $H_{1}$ under several assumption on the neural network class, but it differs from ours in two key respects. First, they consider a single stochastic process $\tilde{\mathcal{K}}_{t}$, while we define a family $\mathcal{K}^{*}_{t}(\lambda)$ parameterized by $\lambda$ (see eq.˜3). This parameterization enables the practical test we introduce in Section 3.3 to adaptively tune $\lambda$ and approximate the optimal growth rate. Second, even when $\lambda$ is fixed, our process exhibits faster expected growth under $H_{1}$. Specifically, if we were to adopt their approach for our one-sided test, we would obtain:
@@ 3.3 Practical DP auditing
$\mathcal{K}_{0}^{*}(\lambda_{1}^{0})=1,\quad\mathcal{K}_{t}^{*}(\lambda_{1}^{t})=\mathcal{K}^{*}_{t-1}(\lambda_{1}^{t-1})\times(1+\lambda_{t}[f^{*}(X_{t})-f^{*}(Y_{t})-\tau(\varepsilon,\delta)]),$
it is easy to see that by restricting the range of $\lambda_{t}$ as in the proof of Theorem˜3.2, $\mathcal{K}_{t}^{*}(\lambda_{1}^{t})$ is a nonnegative supermartingale so long as $\lambda_{t}$ is predictable (i.e, choosen before observing $X_{t},Y_{t}$), implying that, with high probability, $\mathcal{K}^{*}_{t}(\lambda_{1}^{t})$ remains bounded over time under $H_{0}$. Next, we need to choose a predictable sequence $\{\lambda_{t}\}_{t\geq 1}$ that ensures that $\mathcal{K}_{t}(\lambda_{1}^{t})$ grows rapidly under $H_{1}$. This can be done by running ONS with losses $\ell^{*}_{t}(\lambda)=-\log(1+\lambda[f^{*}(X_{t})-f^{*}(Y_{t})-\tau(\varepsilon,\delta)])$. The regret bound of ONS from hazan2007ons , summarized in Theorem˜E.1, implies that for any $\lambda$,
$\sum_{i\in[t]}\ell^{*}_{t}(\lambda)-\sum_{i\in[t]}\ell^{*}_{t}(\lambda_{t})\leq O(\log(t)).$
Noticing that $\sum_{i\in[t]}\ell^{*}_{t}(\lambda)=\log(\mathcal{K}^{*}_{t}(\lambda))$ and $\sum_{i\in[t]}\ell^{*}_{t}(\lambda_{t})=\log(\mathcal{K}^{*}_{t}(\lambda_{1}^{t}))$, we have that for sufficiently large $t$
$\Theta(t\Delta^{2})-\log(\mathcal{K}^{*}_{t}(\lambda_{1}^{t}))\leq\log(\mathcal{K}^{*}_{t}(\lambda^{*}))-\log(\mathcal{K}^{*}_{t}(\lambda_{1}^{t}))\leq O(\log(t)),$
which implies that $\lim_{t\to\infty}\frac{\log(\mathcal{K}^{*}_{t}(\lambda_{1}^{t}))}{t}=\Omega(\Delta^{2})$. Hence, learning $\lambda^{*}$ with ONS does not hurt asymptotically if we know $f^{*}$. Finally, $f^{*}$ can be learned by running OGA with losses $h_{t}(f)=\langle f,K(X_{t},\cdot)-K(Y_{t},\cdot)\rangle_{\mathcal{H}}$:
@@
10:   Send $h_{t}(f)=\langle f,K(X_{t},\cdot)-K(Y_{t},\cdot)\rangle_{\mathcal{H}}$ to OGA, receive $f_{t+1}$
11:   Send $\ell_{t}(\lambda)=-\log(1+\lambda\left[\langle f_{t},K(X_{t},\cdot)-K(Y_{t},\cdot)\rangle_{\mathcal{H}}-\tau(\varepsilon,\delta)\right])$ to ONS, receive $\lambda_{t+1}$
@@ 3.4 An alternative approach based on e-processes
For our concrete problem, we can prove that for every $t\geq 1$, $E_{t}:=\frac{2+f_{t}(X_{t})-f_{t}(Y_{t})}{2+\tau}$ is an e-value for the null if $\{f_{t}\}_{t\geq 1}$ are predictable. This allows to re-write the processes $\mathcal{K}_{t}(\lambda_{1}^{t})$ from the previous sections in the form of eq.˜4, and noting that the optimization of $\{\lambda_{i}\}_{i\geq 1}$ is equivalent to the optimization of $\{\beta_{i}\}_{i\geq 1}$, we are able to design a slightly different sequential DP auditing procedure, Algorithm 2. The details of the algorithm, its statistical properties, and an experimental comparison with Algorithm 1 are provided in Appendix C.
@@ 4.1 Additive-Noise Mechanisms for Mean Estimation
We use the sequential DP tester to evaluate the proposed mechanisms for $\varepsilon\in\{0.01,0.1\}$, controlling the additive noise introduced by each mechanism. For each setting, we test the null hypothesis that the mechanism satisfies ($\varepsilon,\delta$)-DP using the characterization in Definition˜3.2 against the alternative that it does not. For this set of experiments, we fix the neighboring datasets to $S=\{0\}$ and $S^{\prime}=\{0,1\}$, although the sequential test remains agnostic of the specific choice of neighboring datasets. Moreover, we use 20 initial samples to set the bandwidth for the MMD tester using the median of the pairwise distances garreau2017large , which are then excluded from the actual testing phase to maintain statistical validity. We repeat each experiment 20 times and report the aggregated findings to ensure robust results and account for statistical variability. We report a failure to reject the null (no violation detected) when the test reaches 2000 observations for $\varepsilon=0.01$ and 5000 samples for $\varepsilon=0.1$.
@@ 4.2 DP-SGD Auditing in Less-Than-One Training Run
In contrast, for the non-private implementation (Figure˜1(b)), the audit successfully detects violations. It rejects the hypothesis $H_{0}:\varepsilon_{canary}\leq 0.01$ in an average of just 60 observations and $H_{0}:\varepsilon_{canary}=0.1$ in 75 observations. After 250 observations, our method establishes a privacy lower bound of $\varepsilon=0.43$ for this non-private mechanism, and $\varepsilon=0.59$ after 2500 observations (in Appendix˜G, Figure˜7).
@@ 5 Discussion
A first limitation of our approach, as observed for some DP-SGD audits, is the very large sample complexity required for large $\varepsilon$ privacy regimes. This is an inherent limitation of $\mathrm{MMD}$-based auditing that also affects previous batch tests relying on this statistic. A second limitation, shared with prior work, is the reliance on a fixed pair of adjacent datasets. Ideally, we would like to combine our algorithm with a procedure that finds a worst-case pair.
@@
We thank Jamie Hayes for useful guidance on auditing DP-SGD. We thank the anonymous reviewer who pointed out the connection between $\mathrm{MMD}$ and Total Variation, which helped us improve the $\mathrm{MMD}$ bound in Theorem˜3.1. AR is supported by NSF grants DMS-2310718 and IIS-2229881.
@@ Appendix A Improvement on Lower-Bounding Hockey-Stick Divergence with the MMD
Theorem 4.13 in [26] introduces a lower bound on the MMD between two distributions that is parameterized by the order and magnitude of corresponding Hockey-Stick divergence. This connection divergence enables DP auditing with the MMD. Theorem 3.1 below tightens this bound . For completeness, we present first Theorem 4.13 in [26] and then a proof of Theorem˜3.1.
@@
([26, Theorem 4.13]) Let $\mathcal{H}$ be a reproducing kernel Hilbert space with domain $\mathcal{X}$ and kernel $K(\cdot,\cdot)$. Suppose $K(x,x)\leq 1$ for all $x\in\mathcal{X}$. If mechanism $\mathcal{A}$ is $(\varepsilon,\delta)$-DP, then for any $S\sim S^{\prime}$
@@
since $\inf_{\pi\in\Pi(P,Q)}\mathbb{P}_{(X,Y)\sim\pi}[X\neq Y]=\operatorname{TV}(P,Q)$, where $\Pi(P,Q)$ denotes the set of all couplings between $P$ and $Q$ (see, e.g., [27, Proposition 4.7]). Finally, we use that
@@
as stated in [25, Remark A.1.b].
@@
The $\mathrm{MMD}$ bound that we provide in Theorem˜3.1 strictly improves over the one from [26, Theorem 4.13] for all privacy parameters $\varepsilon>0,\delta\in(0,1)$:
@@ Appendix B Proofs in Section˜3
    and the optimization domain is an interval of length $1/(4+2\tau(\varepsilon,\delta))$. Noticing that $\sum_{i\in[t]}\ell_{t}(\lambda)=\log(\mathcal{K}_{t}(\lambda))$ and $\sum_{i\in[t]}\ell_{t}(\lambda_{t})=\log(\mathcal{K}_{t})$, we have that, for sufficiently large $t$,
@@
    Lemma 3 in [44] states that for any $a>0$ and $b\in(0,4]$,
@@
    $\leq\sum_{t\geq 1}\mathbb{P}\left[\mathcal{K}_{t}<1/\alpha\text{ and }\frac{1}{t}\sum_{i\in[t]}[f_{i}(X_{i})-f(Y_{i})-\tau(\varepsilon,\delta)]\geq\Delta/2\right]$
@@
    where $t_{1}=\min\{n\in\mathbb{N}:\exp\left(\frac{t\Delta^{2}}{768}-10\log(t)\right)\geq 1/\alpha\}$, since $\frac{t\Delta^{2}}{768}-10\log(t)$ is increasing in $t$. Finally, if $\log(t)/t\leq\frac{\Delta^{2}/2}{7680}$, then $\frac{t\Delta^{2}}{768}-10\log(t)\geq\frac{t\Delta^{2}/2}{768}$. Lemma 3 from [44] implies that $\log(t)/t\leq\frac{\Delta^{2}/2}{7680}$ for $t=\Omega(\frac{\log(1/\Delta^{2})}{\Delta^{2}})$. In addition, it is easy to see that $\frac{t\Delta^{2}/2}{768}\geq\log(1/\alpha)$ for $t=\Omega\big{(}\frac{\log(1/\alpha)}{\Delta^{2}}\big{)}$. We conclude that $t=\Omega\big{(}\frac{\log(1/\Delta^{2})+\log(1/\alpha)}{\Delta^{2}}\big{)}$ suffices to obtain

### v2

@@ 1 Introduction
Our specific test statistics are built upon the Maximum Mean Discrepancy (MMD) gretton2012kernel, a powerful kernel-based metric for comparing probability distributions. MMD is particularly well-suited for the two-sample tests emerging in privacy auditing, as bounds on MMD can be easily translated to approximate DP parameters 2024DPAuditorium. Further, MMD is highly flexible. It can incorporate prior knowledge in white-box settings—with access to intermediate gradients or the noise distribution family—or, as our experiments show, operate effectively in a black-box manner.
@@
  • We introduce a new general MMD-based one-sided sequential testing framework. As a key application, we instantiate this framework for auditing approximate DP mechanisms (see Section˜3). Our method enables anytime valid inference while controlling Type I error, eliminating the need for pre-defined test sample sizes, and ensuring a bounded expected stopping time under the alternative. We establish these theoretical guarantees in Theorems 3.2 and 3.3.
@@
  • We validate our methods on common DP mechanisms with Gaussian and Laplace noise. We further demonstrate efficacy on auditing benchmark algorithms (2024DPAuditorium; dpsniper) and provide results for the challenging case of DP-SGD abadi2016deep; song2013stochastic, showcasing the practical benefits of early failure detection enabled by our sequential approach.
@@ 3 Sequential Auditing
This section formalizes the proposed auditing by betting framework. We start by formally introducing the problem statement and continue by introducing an MMD-based test statistic. Subsequently, we develop Algorithm 1, a sequential test for privacy auditing. We then provide formal guarantees on controlling Type I error uniformly over the null, and demonstrating exponential growth under the alternative, which implies a finite stopping time when detecting faulty algorithms. We conclude this section with a comparative discussion of two subroutines within the main sequential algorithm.
@@ 3.1 Problem statement
Theorem˜3.1 offers a strict improvement over (2024DPAuditorium, Theorem 4.13). Further, the new bound has the nice property of not becoming vacuous as $\varepsilon$ increases. In fact, it approaches $\sqrt{2}$ as $\varepsilon\to\infty$, while the previous bound grows to infinity with $e^{\varepsilon}$.
@@ 3.2 An abstract template for $(\varepsilon,\delta)$-DP sequential testing
$\mathcal{K}_{0}^{*}(\lambda)=1,\quad\mathcal{K}_{t}^{*}(\lambda)=\mathcal{K}^{*}_{t-1}(\lambda)\times(1+\lambda[f^{*}(X_{t})-f^{*}(Y_{t})-\tau(\varepsilon,\delta)]),$ (3)
@@
We are not the first to draw inspiration from shekhar2023nonparametric2sampletesting to design a one-sided test. Recently, richter2024auditing proposed a one-sided test in the context of detecting behavioral shifts in language models for the hypotheses $H_{0}:D(\mathbb{P},\mathbb{Q})\leq\tau$ versus $H_{1}:D(\mathbb{P},\mathbb{Q})>\tau$ for some $\tau>0$, where the distance function is defined as $D(\mathbb{P},\mathbb{Q})=\sup_{f\in\mathcal{F}}\left|\mathbb{E}_{X\sim\mathbb{P},\,Y\sim\mathbb{Q}}[f(X)-f(Y)]\right|$ and $\mathcal{F}$ is a class of neural networks with scalar outputs in [1/2,1/2][-1/2,1/2]. Their method also constructs a stochastic process that is a supermartingale under $H_{0}$ and can grow under $H_{1}$ under several assumption on the neural network class, but it differs from ours in two key respects. First, they consider a single stochastic process $\tilde{\mathcal{K}}_{t}$, while we define a family $\mathcal{K}^{*}_{t}(\lambda)$ parameterized by $\lambda$ (see eq.˜3). This parameterization enables the practical test we introduce in Section 3.3 to adaptively tune $\lambda$ and approximate the optimal growth rate. Second, even when $\lambda$ is fixed, our process exhibits faster expected growth under $H_{1}$. Specifically, if we were to adopt their approach for our one-sided test, we would obtain: [^1] Their process does not include the factors of $1/2$ from our expression; we include them to standardize both settings, as we work with function classes outputting in [1,1][-1,1] whereas theirs output in [1/2,1/2][-1/2,1/2].
@@ 3.3 Practical DP auditing
$\mathcal{K}_{0}^{*}=1,\quad\mathcal{K}_{t}^{*}=\mathcal{K}^{*}_{t-1}\times(1+\lambda_{t}[f^{*}(X_{t})-f^{*}(Y_{t})-\tau(\varepsilon,\delta)]),$
it is easy to see that by restricting the range of $\lambda_{t}$ as in the proof of Theorem˜3.2, $\mathcal{K}_{t}^{*}$ is a nonnegative supermartingale so long as $\lambda_{t}$ is predictable (i.e, choosen before observing $X_{t},Y_{t}$), implying that, with high probability, $\mathcal{K}^{*}_{t}$ remains bounded over time under $H_{0}$. Next, we need to choose a predictable sequence $\{\lambda_{t}\}_{t\geq 1}$ that ensures that $\mathcal{K}_{t}$ grows rapidly under $H_{1}$. This can be done by running ONS with losses $\ell^{*}_{t}(\lambda)=-\log(1+\lambda[f^{*}(X_{t})-f^{*}(Y_{t})-\tau(\varepsilon,\delta)])$. The regret bound of ONS from hazan2007ons, summarized in Theorem˜E.1, implies that for any $\lambda$,
$\sum_{i\in[t]}\ell^{*}_{i}(\lambda)-\sum_{i\in[t]}\ell^{*}_{i}(\lambda_{i})\leq O(\log(t)).$
Noticing that $\sum_{i\in[t]}\ell^{*}_{i}(\lambda)=\log(\mathcal{K}^{*}_{t}(\lambda))$, where $\log(\mathcal{K}^{*}_{t}(\lambda))$ is the process from Equation˜3 and $\sum_{i\in[t]}\ell^{*}_{i}(\lambda_{i})=\log(\mathcal{K}^{*}_{t})$, we have that
$\lim_{t\to\infty}\frac{\log(\mathcal{K}^{*}_{t})}{t}=\lim_{t\to\infty}\frac{\log(\mathcal{K}^{*}_{t}(\lambda^{*}))}{t}=\Omega(\Delta^{2}),$
where the last equality comes from Theorem˜3.2, Part 2. Hence, learning $\lambda^{*}$ with ONS does not hurt asymptotically if we know $f^{*}$. Finally, $f^{*}$ can be learned by running OGA with losses $h_{t}(f)=\langle f,K(X_{t},\cdot)-K(Y_{t},\cdot)\rangle_{\mathcal{H}}$:
@@
10:   Send $h_{t}(f_{t})=\langle f_{t},K(X_{t},\cdot)-K(Y_{t},\cdot)\rangle_{\mathcal{H}}$ to OGA, receive $f_{t+1}$
11:   Send $\ell_{t}(\lambda_{t})=-\log(1+\lambda_{t}\left[\langle f_{t},K(X_{t},\cdot)-K(Y_{t},\cdot)\rangle_{\mathcal{H}}-\tau(\varepsilon,\delta)\right])$ to ONS, get $\lambda_{t+1}$
@@ 3.4 An alternative approach based on e-processes
For our concrete problem, we can prove that for every $t\geq 1$, $E_{t}:=\frac{2+f_{t}(X_{t})-f_{t}(Y_{t})}{2+\tau}$ is an e-value for the null if $\{f_{t}\}_{t\geq 1}$ are predictable. This allows to re-write the processes $\mathcal{K}_{t}$ from the previous sections in the form of eq.˜4, and noting that the optimization of $\{\lambda_{i}\}_{i\geq 1}$ is equivalent to the optimization of $\{\beta_{i}\}_{i\geq 1}$, we are able to design a slightly different sequential DP auditing procedure, Algorithm 2. The details of the algorithm, its statistical properties, and an experimental comparison with Algorithm 1 are provided in Appendix C.
@@ 4.1 Additive-Noise Mechanisms for Mean Estimation
We use the sequential DP tester to evaluate the proposed mechanisms for $\varepsilon\in\{0.01,0.1\}$, controlling the additive noise introduced by each mechanism. For each setting, we test the null hypothesis that the mechanism satisfies ($\varepsilon,\delta$)-DP using the characterization in Definition˜3.2 against the alternative that it does not. For this set of experiments, we fix the neighboring datasets to $S=\{0\}$ and $S^{\prime}=\{0,1\}$, although the sequential test remains agnostic of the specific choice of neighboring datasets. Moreover, we use 20 initial samples to set the bandwidth for the MMD tester using the median of the pairwise distances garreau2017large, which are then excluded from the actual testing phase to maintain statistical validity. We repeat each experiment 20 times and report the aggregated findings to ensure robust results and account for statistical variability. We report a failure to reject the null (no violation detected) when the test reaches 2,000 observations for $\varepsilon=0.01$ and 5,000 samples for $\varepsilon=0.1$.
@@ 4.2 DP-SGD Auditing in Less-Than-One Training Run
In contrast, for the non-private implementation (Figure˜1(b)), the audit successfully detects violations. It rejects the hypothesis $H_{0}:\varepsilon_{canary}\leq 0.01$ in an average of just 60 observations and $H_{0}:\varepsilon_{canary}=0.1$ in 75 observations. After 250 observations, our method establishes a privacy lower bound of $\varepsilon=0.43$ for this non-private mechanism, and $\varepsilon=0.59$ after 2,500 observations (in Appendix˜G, Figure˜7).
@@ 5 Discussion
A first limitation of our approach, as observed for some DP-SGD audits, is the very large sample complexity required for large $\varepsilon$ privacy regimes. This is an inherent limitation of $\mathrm{MMD}$-based auditing that also affects previous batch tests relying on this statistic. A second limitation, shared with prior work, is the reliance on a fixed pair of adjacent datasets. Ideally, we would like to combine our algorithm with a procedure that adaptively finds a worst-case pair.
@@
We thank Jamie Hayes for useful guidance on auditing DP-SGD. We thank the anonymous reviewer who pointed out the connection between $\mathrm{MMD}$ and Total Variation, which helped us improve the $\mathrm{MMD}$ bound in Theorem˜3.1. AR is supported by NSF grants DMS-2310718 and IIS-2229881. TG was partially funded to attend this conference by the CMU GSA/Provost Conference Funding.
@@ Appendix A Improvement on Lower-Bounding Hockey-Stick Divergence with the MMD
Theorem 4.13 in [2024DPAuditorium] introduces a lower bound on the MMD between two distributions that is parameterized by the order and magnitude of corresponding Hockey-Stick divergence. This connection divergence enables DP auditing with the MMD. Theorem 3.1 below tightens this bound . For completeness, we present first Theorem 4.13 in [2024DPAuditorium] and then a proof of Theorem˜3.1.
@@
([2024DPAuditorium, Theorem 4.13]) Let $\mathcal{H}$ be a reproducing kernel Hilbert space with domain $\mathcal{X}$ and kernel $K(\cdot,\cdot)$. Suppose $K(x,x)\leq 1$ for all $x\in\mathcal{X}$. If mechanism $\mathcal{A}$ is $(\varepsilon,\delta)$-DP, then for any $S\sim S^{\prime}$
@@
since $\inf_{\pi\in\Pi(P,Q)}\mathbb{P}_{(X,Y)\sim\pi}[X\neq Y]=\operatorname{TV}(P,Q)$, where $\Pi(P,Q)$ denotes the set of all couplings between $P$ and $Q$ (see, e.g., [levin2017MarkovChainsAndMixing, Proposition 4.7]). Finally, we use that
@@
as stated in [kairouz2015composition_of_DP, Remark A.1.b].
@@
The $\mathrm{MMD}$ bound that we provide in Theorem˜3.1 strictly improves over the one from [2024DPAuditorium, Theorem 4.13] for all privacy parameters $\varepsilon>0,\delta\in(0,1)$:
@@ Appendix B Proofs in Section˜3
    and the optimization domain is an interval of length $1/(4+2\tau(\varepsilon,\delta))$. Noticing that $\sum_{i\in[t]}\ell_{i}(\lambda)=\log(\mathcal{K}_{t}(\lambda))$ and $\sum_{i\in[t]}\ell_{i}(\lambda_{i})=\log(\mathcal{K}_{t})$, we have that, for sufficiently large $t$,
@@
    Lemma 3 in [shekhar2023nonparametric2sampletesting] states that for any $a>0$ and $b\in(0,4]$,
@@
    $\leq\sum_{t\geq 1}\mathbb{P}\left[\mathcal{K}_{t}<1/\alpha\text{ and }\frac{1}{t}\sum_{i\in[t]}[f_{i}(X_{i})-f_{i}(Y_{i})-\tau(\varepsilon,\delta)]\geq\Delta/2\right]$
@@
    where $t_{1}=\min\{n\in\mathbb{N}:\exp\left(\frac{t\Delta^{2}}{768}-10\log(t)\right)\geq 1/\alpha\}$, since $\frac{t\Delta^{2}}{768}-10\log(t)$ is increasing in $t$. Finally, if $\log(t)/t\leq\frac{\Delta^{2}/2}{7680}$, then $\frac{t\Delta^{2}}{768}-10\log(t)\geq\frac{t\Delta^{2}/2}{768}$. Lemma 3 from [shekhar2023nonparametric2sampletesting] implies that $\log(t)/t\leq\frac{\Delta^{2}/2}{7680}$ for $t=\Omega(\frac{\log(1/\Delta^{2})}{\Delta^{2}})$. In addition, it is easy to see that $\frac{t\Delta^{2}/2}{768}\geq\log(1/\alpha)$ for $t=\Omega\big(\frac{\log(1/\alpha)}{\Delta^{2}}\big)$. We conclude that $t=\Omega\big(\frac{\log(1/\Delta^{2})+\log(1/\alpha)}{\Delta^{2}}\big)$ suffices to obtain